You bank statement shows a current balance of AED 100,000 as a result of a single deposit made 10 years previously. Over this period the interest rate has been 5%. The original deposit is nearest to

AED 39,000

AED 42,000

AED 32,550

AED 48,550
Answer #1

Use future value rule as FV = PV(1 + r)^n

Here FV is 100000, r is 5% and n is 10 years

Find original deposit PV as

100000 = PV(1 + 5%)^10

PV = 100000*(1 + 5%)^-10 = 61391

Closest value is AED 48,550